i have hti sproblem x^2+8x-y+6=0 i need to find the focus,directrix and the vertex. where i get lost at is -y+6=-x^2-8x then hwat is next. thank you for your help
Answer by stanbon(48569)

Write that as y-6= x^2+8x
Now, complete the square on the right side to get:
x^2+8x+16 = y-6+16
(x+4)^2 = y+10
Vertex is at (-4,-10)
The parabola opens up so directrix is 1/4 below vertex
at y = -41/4
Focus is at 1/4 above vertex at (-4,-39/4)
Cheers,
Stan H.